# Scaling Laws for Precision

### v1

@@ 1 Introduction
Scale has emerged as a central driver of progress in deep learning (Brown, 2020). Key work on scaling (Kaplan et al., 2020; Hoffmann et al., 2022) studied tradeoffs between model/dataset size to balance performance and compute. However, the precision in which models are trained and served is an important third factor that contributes to both cost and performance. Deep learning is trending towards lower precision: current frontier models like the Llama-3 series are trained in BF16 (Dubey et al., 2024), and there is widespread effort to move the pretraining paradigm to FP8 (Micikevicius et al., 2022). The next generation of hardware will support FP4, and advances in weight-only quantization have led to training in binary and ternary at scale (Ma et al., 2024; Wang et al., 2023). How far will these paradigms go? Specifically, we ask:
@@
Figure 1: Schematic of key findings. (Left) Training a fixed model size to various data budgets in BF16 and quantizing weights at the end. We find that degradation due to post-train quantization increases with tokens seen during pretraining, so that eventually additional pretraining data can be harmful. (Right) Our scaling suggests training larger models in lower precision can be compute-optimal according to the cost model in Section 4.3. Weights, activations, attention quantized, all models trained on the same data budget, details in Appendix H.
@@ 2.1 Quantization Fundamentals: How, What, When
Quantization: How. Quantization of an operation typically refers to rounding of values in matrices involved in some computation on the forward/backward pass, with accumulation of gradients in high/full precision. Quantization is usually done to integer or floating-point type.
Quantization: What. Only weights. “Quantization-aware training” Quantizing only weights during training does not offer any compute savings because matrix multiplications are still done in high precision. However, this is commonly done to allow weights to adapt to low precision so they can be served at very low precision at inference-time, thereby alleviating memory bottlenecks (Ma et al., 2024; Wang et al., 2023). We will refer to this as “quantization-aware-training.”
Weights, activations, attention. “Low-precision training” Quantizing and activations and attention in addition to weights allows for compute gains because matrix multiplications can be done in low precision (if the hardware supports it) since everything is in the same precision. FP8 training on the Hopper line of GPUs is an example (Micikevicius et al., 2022). We will refer to this setting as “low-precision training” to distinguish it from quantization-aware training.
Quantization: When. Quantization can be done during or after training. In practice, when seeking to reduce inference-time memory costs, one first attempts post-train quantization. If that degrades the model too much, quantization-aware-training is used. Post-train quantization is typically only applied to model weights (Frantar et al., 2022; Dettmers et al., 2022; Lin et al., 2023; Xiao et al., 2023). To reduce pretraining costs, low-precision-training is needed. We will study scaling laws for post-training quantization in Section 3, for quantized training in Section 4 (examining both quantization-aware training and low precision training) and unify the two in Section 5. The numerical values of all our fitted constants can be found in Appendix I.
@@ 3.1 Overtrained Models Degrade more when Post-Train Quantized
where $C_{T},\gamma_{D},\gamma_{N},\gamma_{\text{post}}$ are positive fitted constants. As we find the fitted values of $\gamma_{D}$ and $\gamma_{N}$ to be similar (see Appendix I for numerical values), we can think of this as an approximate power law in the token/parameter ratio $D/N$. The intuition for this poor data scaling might be that as models train on more data, they compress more information into their weights, so that perturbations to weights in the form of quantization are more harmful to loss, all else equal. We discuss formal theoretical interpretations in Appendix G.
This finding implies that for models that will be post-train quantized, there exists an amount of pretraining data beyond which additional data is actively harmful to performance at inference-time (see top-left, Figure 2). This can be defined as the point where additional data increases post-train degradation more than it decreases loss during pretraining. We solve analytically for this critical data size in Appendix E. We thus summarize our first scaling finding as follows.
@@ 4 Scaling Laws for Quantized Training
In this section we study pretraining with weights, activations, and KV cache in various precisions. Importantly, only training precision, not test-time precision, is varied in this section; we discuss the interaction between train and test-time precision in Section 5. We sweep the training precisions of the weights, activations, and KV cache $P_{\text{w}},P_{\text{a}},P_{\text{kv}}\in[3,12]$ individually, as well as training BF16 baselines. We also pretrain models with arbitrary combinations of $P_{\text{w}},P_{\text{a}},P_{\text{kv}}$ to validate our scaling laws. To perform quantization during training, we follow the standard specification of Micikevicius et al. (2022) unless otherwise noted (see Appendix D for implementation details).
@@
In line with the empirical trends in Figure 3, we find the best fit for the tradeoff between weight precision and parameters is $N_{\text{eff}}(N,P_{\text{w}})=N(1-e^{-P_{\text{w}}/\gamma_{\text{w}}})$, where $\gamma_{\text{w}}$ is a fitted constant measuring the sensitivity of model weights (alternative fits explored in Appendix I). We therefore modify Chinchilla scaling to account for $N_{\text{eff}}$ by making the substitution $N\mapsto N_{\text{eff}}(N,P_{\text{w}})$, giving the modified form:
@@
Quantization-aware training does not change the cost of pretraining. This is because modern GPUs require inputs to a matrix multiplication to have the same precision, i.e. $P_{\text{w}}=P_{\text{a}}=P_{\text{kv}}$ (Micikevicius et al., 2022). To understand the interplay between precision and pretraining compute we must now analyze the scaling behavior of $P_{\text{a}}$ and $P_{\text{kv}}$ as well.
@@
We test whether this independence approximately holds by comparing the predictive power of a model with marginally fitted constants and a model with jointly fitted constants. We show the predictive power of both models in Figure 5(b, c), finding that both methods for fitting constants have approximately the same predictive power. These results suggest that the independence assumption is reasonable. We both present further evidence that this “factorized” functional form is a strong fit to the data as well as discuss alternative factorization schemes in Appendix K.
@@ 4.3 Implications For Pretraining
When training in a precision $P$, meaning $P_{\text{w}}=P_{\text{a}}=P_{\text{kv}}=P$, compute cost scales linearly in $P$ (Abdelkhalik et al., 2022; Peng et al., 2023) [^2] In practice, the gains are less than linear due to systems overhead.. Hoffmann et al. (2022) performed all experiments in 16-bit precision and use a cost model of $C=6ND$ FLOPs. We generalize this to $C=\frac{6}{16}NDP$ to account for the linear relation between compute and precision, which reduces to the Chinchilla cost function for $P=16$. We now examine three practically relevant variants of the following optimization problem.
@@ 4.3.2 Compute-Optimal Pretraining Precision is in General Independent of Compute
which reveals that in general the optimal pretraining precision is independent of compute budget. This suggests that compute-optimal precision should be held fixed to $P^{*}$ while $N,D$ are scaled according to Equation 6. We find this $P^{*}\approx 7$ bits when fitting our scaling law on runs with quantization done to integer type. This has two consequences: first, this means the de-facto practice of training models in BF16 may be suboptimal. Second, the race to low-precision training will likely have to stop before going below 4-bits, since this would force model sizes to become disproportionately (more than 4x) larger to maintain loss scaling (see Figure 3, left).
We test our predictions in Figure 6 at a larger scale. We train compute-matched models at various parameter count and precision ranging from FP4 to FP32 and 220M to 1.6B parameters. We train in floating-point type since that is standard in pretraining (Groeneveld et al., 2024; Deitke et al., 2024), though our scaling laws are fitted on integer type. We plot our predicted trend in Figure 6 (left) and the empirical values in the middle. We find that scaling fits on integer type are a strong fit until 4-bit precision, at which points the difference between the two types becomes more apparent. Integer fits assume all bits contribute in the same way, but the split of floating-point bits into exponent and mantissa means that each likely has its own scaling behavior. The matching of qualitative trends throughout, with the optimum being close to the predicted optimum of $P^{*}\approx 7$ bits suggests that similar scaling laws may exist across types.
@@ 5 A Unified Scaling Law for Precision
Two competing effects at play during post-train quantization. Intuitively, training any of $P_{\text{w}},P_{\text{a}},P_{\text{kv}}$ in low precision forces the model to learn weights that are robust to “quantization noise,” so they degrade less under PTQ. However, the reduced $N\mapsto N_{\text{eff}}$ effective parameter count of a model trained in low precision suggests that models trained in low precision will degrade more because $\delta_{\text{PTQ}}$ increases with $N^{-\gamma_{N}}$ as we found in Section 3. We call this second effect the “overtraining” effect. In practice, the first “robustification” effect wins out, so that models trained in lower precision overall degrade less when post-train quantized compared to models trained in high precision. We confirm using $N_{\text{eff}}$ rather than $N$ to predict degradation given various training precisions leads to a substantially stronger fit in Figure 16(top left, top center), to verify the competing overtraining effect is real.
@@ 6 Conclusion and Limitations
We find that the common inference-time technique of post-train quantization can incur large degradation at very high data budgets, demonstrating a striking example of how more pretraining compute does not always imply stronger models at inference-time. Seeking better data scaling, we study quantization-aware and low precision training. We find that parameters and bit precision are well modeled as interchangeably controlling an “effective parameter count” of the model allows us to predict finite-precision loss effects accurately during both training and inference. The resulting scaling law makes surprising predictions that we qualitatively validate on language models pretrained from scratch with up to 1.7B parameters.
@@ 7 Acknowledgements
Tanishq Kumar thanks Tim Dettmers, Chris De Sa, Neil Band and Luke Bailey for helpful comments and discussion. Blake Bordelon is supported by a Google PhD Fellowship. Cengiz Pehlevan is supported by NSF grant DMS-2134157, NSF CAREER Award IIS-2239780, and a Sloan Research Fellowship. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. Aditi Raghunathan acknowledges support from AI2050 program by Schmidt Sciences (Grant G2264481), Google Research Scholar, Apple, NSF, Cisco.
@@ Appendix A Hyperparameter Details and Ablations
We launch over 20 runs for each $(N,D)$ combination to study scaling in precision, trained and validated on the common crawl split of the Dolma dataset [Soldaini et al., 2024]. We use a Transformer++ implementation: SwiGLU activations [Shazeer, 2020], RoPE embeddings [Su et al., 2021], RMSLayerNorm, Adam $\beta$ values of $(0.9,0.95)$. We adopt a cosine learning rate schedule with 10% warmup period and peak learning rate of 6e-4 for the smallest model and learning rates scaled with width and depth according to depth-$\mu$P for the larger models [Yang et al., 2022, Bordelon et al., 2023]. We use a sequence length of 1024 and batch size of 256 throughout, with Adam $\epsilon$ 1e-15, following [Wortsman et al., 2023a]. We use weight decay of $0.1$, as [Ahmadian et al., 2023] find some results in the quantization literature may be artifacts of insufficient weight decay. We follow [Ma et al., 2024] in including a LayerNorm before projections because they find it is important for low precision training to be stable. These are the hyperparameters and settings used for the main scaling law experiments.
@@ Appendix B Additional Related Work
Low precision has been key to improving the efficiency of training and using LLMs [Micikevicius et al., 2017, Shoeybi et al., 2019, Wortsman et al., 2023b, Zhu et al., 2023]. Prior works generally study either precision during training [Courbariaux et al., 2014, Dettmers et al., 2024, 2021, Sun et al., 2020, Liu et al., 2023] or the effects of changing the precision after training (post-training quantization) [Frantar et al., 2022, Lin et al., 2024b, Dettmers et al., 2022, Xiao et al., 2023, Sheng et al., 2023, Dettmers et al., 2023]. In this work we study both, the precision during training and after, and unify them from a scaling perspective.
@@
By scaling up the transformer architecture [Vaswani et al., 2017] a variety of large language models have been proposed [Brown, 2020, Rae et al., 2021, Touvron et al., 2023a, b, Dubey et al., 2024, Le Scao et al., 2023, Muennighoff et al., 2022, 2024a, Groeneveld et al., 2024, Jiang et al., 2023, Zhang et al., 2022, Allal et al., 2023, Li et al., 2023, Lozhkov et al., 2024, Luukkonen et al., 2023, Bai et al., 2023, Chowdhery et al., 2023, Team et al., 2023, Üstün et al., 2024, Deitke et al., 2024]. To improve our understanding of these models, various works have investigated their scaling properties [Ruan et al., 2024, Allen-Zhu and Li, 2024, Hägele et al., 2024]. Many aspects are relevant to scaling including the architecture [Tay et al., 2022a, Krajewski et al., 2024, Tao et al., 2024, Clark et al., 2022, Tay et al., 2022b, Scao et al., 2022, Peng et al., 2024], the modalities considered [Aghajanyan et al., 2023, Alabdulmohsin et al., 2022, Cherti et al., 2023], the performance metrics [Wei et al., 2022, Srivastava et al., 2022, Isik et al., 2024], the data composition [Li et al., 2024, Liu et al., 2024, Albalak et al., 2024] and data repetitions [Muennighoff et al., 2024b]. Our work analyzes one such aspect, which is key to better scaling: the numeric precision during and after training.

### v2

@@ 1 Introduction
Scale has emerged as a central driver of progress in deep learning (Brown et al., 2020). Key work on scaling (Kaplan et al., 2020; Hoffmann et al., 2022) studied tradeoffs between model/dataset size to balance performance and compute. However, the precision in which models are trained and served is an important third factor that contributes to both cost and performance. Deep learning is trending towards lower precision: current frontier models like the Llama-3 series are trained in BF16 (Dubey et al., 2024), and there is widespread effort to move the pretraining paradigm to FP8 (Micikevicius et al., 2022). The next generation of hardware will support FP4, and advances in weight-only quantization have led to training in binary and ternary at scale (Ma et al., 2024; Wang et al., 2023). How far will these paradigms go? Specifically, we ask:
@@
Figure 1: Schematic of key findings. (Left) Training a fixed model size to various data budgets in BF16 and quantizing weights at the end. We find that degradation due to post-train quantization increases with tokens seen during pretraining, so that eventually additional pretraining data can be harmful. (Right) Our scaling suggests training larger models in lower precision can be compute-optimal according to the cost model in Section 4.3. Weights, activations, attention quantized, all models trained on the same data budget, details in Appendix J.
@@ 2.1 Quantization Fundamentals: How, What, When
Quantization: How. Quantization of an operation typically refers to rounding of values in matrices involved in some computation on the forward or backward pass, depending on what is quantized, and when. Quantization is usually done to integer or floating-point type.
Quantization: What. Only weights. “Quantization-aware training” Quantizing only weights during training does not offer any compute savings because matrix multiplications are still done in high precision. However, this is commonly done to allow weights to adapt to low precision so they can be served at very low precision at inference-time, thereby alleviating memory bottlenecks (Ma et al., 2024; Wang et al., 2023). We will refer to this as “quantization-aware-training” and defer additional discussion to Appendix D.
Weights, activations, attention. “Low-precision training” Quantizing and activations and attention in addition to weights allows for compute gains because matrix multiplications can be done in low precision (if the hardware supports it) since everything is in the same precision. We will refer to this setting as “low-precision training” to distinguish it from quantization-aware training.
Quantization: When. Quantization can be done during or after training. In practice, when seeking to reduce inference-time memory costs, one first attempts post-train quantization. If that degrades the model too much, quantization-aware-training is used. Post-train quantization is typically only applied to model weights (Frantar et al., 2022; Dettmers et al., 2022; Lin et al., 2023; Xiao et al., 2023). To reduce pretraining costs, low-precision-training is needed. We will study scaling laws for post-training quantization in Section 3, for quantized training in Section 4 (examining both quantization-aware training and low precision training) and unify the two in Section 5. The numerical values of all our fitted constants can be found in Appendix K.
@@ 3.1 Overtrained Models Degrade more when Post-Train Quantized
where $C_{T},\gamma_{D},\gamma_{N},\gamma_{\text{post}}$ are positive fitted constants. As we find the fitted values of $\gamma_{D}$ and $\gamma_{N}$ to be similar (see Appendix K for numerical values), we can think of this as an approximate power law in the token/parameter ratio $D/N$. The intuition for this poor data scaling might be that as models train on more data, they compress more information into their weights, so that perturbations to weights in the form of quantization are more harmful to loss, all else equal. We discuss formal theoretical interpretations in Appendix H.
This finding implies that for models that will be post-train quantized, there exists an amount of pretraining data beyond which additional data is actively harmful to performance at inference-time (see top-left, Figure 2). This can be defined as the point where additional data increases post-train degradation more than it decreases loss during pretraining. We solve analytically for this critical data size in Appendix E, as well analyze a cost model for workloads where inference-cost is the primary concern. We thus summarize our first scaling finding as follows.
@@ 4 Scaling Laws for Quantized Training
In this section we study pretraining with weights, activations, and KV cache in various precisions. Importantly, only training precision, not test-time precision, is varied in this section; we discuss the interaction between train and test-time precision in Section 5. We sweep the training precisions of the weights, activations, and KV cache $P_{\text{w}},P_{\text{a}},P_{\text{kv}}\in[3,12]$ individually, as well as training BF16 baselines. We also pretrain models with arbitrary combinations of $P_{\text{w}},P_{\text{a}},P_{\text{kv}}$ to validate our scaling laws. To perform quantization during training, we quantize the forward pass in integer type unless otherwise noted, see Appendix D for implementation details.
@@
In line with the empirical trends in Figure 3, we find the best fit for the tradeoff between weight precision and parameters is $N_{\text{eff}}(N,P_{\text{w}})=N(1-e^{-P_{\text{w}}/\gamma_{\text{w}}})$, where $\gamma_{\text{w}}$ is a fitted constant measuring the sensitivity of model weights (alternative fits explored in Appendix K). We therefore modify Chinchilla scaling to account for $N_{\text{eff}}$ by making the substitution $N\mapsto N_{\text{eff}}(N,P_{\text{w}})$, giving the modified form:
@@
Quantization-aware training does not change the cost of pretraining. This is because modern GPUs require inputs to a matrix multiplication to have the same precision, i.e. $P_{\text{w}}=P_{\text{a}}=P_{\text{kv}}$ (Micikevicius et al., 2022). To understand the interplay between precision and pretraining compute we must now analyze the scaling behavior of $P_{\text{a}}$ and $P_{\text{kv}}$ as well. Note that in our training experiments, we only quantize on the forward pass to ensure a fair comparison between quantization-aware-training (weights only) and the additional quantization to activations/KV cache, see Appendix D.
@@
We test whether this independence approximately holds by comparing the predictive power of a model with marginally fitted constants and a model with jointly fitted constants. We show the predictive power of both models in Figure 5(b, c), finding that both methods for fitting constants have approximately the same predictive power. These results suggest that the independence assumption is reasonable. We both present further evidence that this “factorized” functional form is a strong fit to the data as well as discuss alternative factorization schemes in Appendix M.
@@ 4.3 Implications For Pretraining
When training in a precision $P$, meaning $P_{\text{w}}=P_{\text{a}}=P_{\text{kv}}=P$, compute cost scales linearly in $P$ (Abdelkhalik et al., 2022) [^2] In practice, the gains are less than linear due to systems overhead.. Hoffmann et al. (2022) performed all experiments in 16-bit precision and use a cost model of $C=6ND$ FLOPs. We generalize this to $C=\frac{6}{16}NDP$ to account for the linear relation between compute and precision, which reduces to the Chinchilla cost function for $P=16$. We now examine three practically relevant variants of the following optimization problem.
@@ 4.3.2 Compute-Optimal Pretraining Precision is in General Independent of Compute
which reveals that in general the optimal pretraining precision is independent of compute budget. This suggests that compute-optimal precision should be held fixed to $P^{*}$ while $N,D$ are scaled according to Equation 6. We find this $P^{*}$ to be around 7-8 bits when fitting our scaling law on runs with quantization done to integer type. This has two consequences: first, this means the de-facto practice of training models in 16-bit may be suboptimal. Second, the race to low-precision training may have to stop before going below 4-bits, since this would force model sizes to become disproportionately (more than 4x) larger to maintain loss scaling (see Figure 3, left).
We test our predictions in Figure 6 at a larger scale. We train compute-matched models at various parameter count and precision ranging from FP4 to FP32 and 220M to 1.6B parameters. We train in floating-point type since that is standard in pretraining (Groeneveld et al., 2024; Deitke et al., 2024), though our scaling laws are fitted on integer type. We plot our predicted trend in Figure 6 (left) and the empirical values in the middle. We find that scaling fits on integer type are a strong fit until 4-bit precision, at which points the difference between the two types becomes more apparent. The matching of qualitative trends throughout, with the optimum being close to the predicted optimum of $P^{*}$ near 7-8 bits suggests that similar scaling laws may exist across types. We initiate a similar analysis for floating-point type in Appendix N.
@@ 5 A Unified Scaling Law for Precision
Two competing effects at play during post-train quantization. Intuitively, training any of $P_{\text{w}},P_{\text{a}},P_{\text{kv}}$ in low precision forces the model to learn weights that are robust to “quantization noise,” so they degrade less under PTQ. But the reduced $N\mapsto N_{\text{eff}}$ implies that models trained in low precision will degrade more because $\delta_{\text{PTQ}}$ increases with $N^{-\gamma_{N}}$ (Section 3). We call this second effect the “overtraining” effect. In practice, the first “robustification” effect is larger, so that models trained in lower precision overall degrade less when post-train quantized. We confirm using $N_{\text{eff}}$ rather than $N$ to predict degradation given various training precisions leads to a substantially stronger fit in Figure 21(top left, top center), to verify the competing overtraining effect.
@@ 6 Conclusion and Limitations
We find that the common inference-time technique of post-train quantization can incur large degradation at very high data budgets, demonstrating a striking example of how more pretraining compute does not always imply stronger models at inference-time. Seeking better data scaling, we study quantization-aware and low precision training. We find that parameters and bit precision are well modeled as interchangeably controlling an “effective parameter count” of the model allows us to predict finite-precision loss effects accurately during both training and inference.
@@ 7 Acknowledgements
Tanishq Kumar thanks Tim Dettmers, Chris De Sa, Neil Band and Luke Bailey for helpful comments and discussion, as well as Ludwig Schmidt for spotting an early typo. Blake Bordelon is supported by a Google PhD Fellowship. Cengiz Pehlevan is supported by NSF grant DMS-2134157, NSF CAREER Award IIS-2239780, and a Sloan Research Fellowship. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. Aditi Raghunathan acknowledges support from AI2050 program by Schmidt Sciences (Grant G2264481), Google Research Scholar, Apple, NSF, Cisco.
@@ Appendix A Hyperparameter Details and Ablations
We launch over 20 runs for each $(N,D)$ combination to study scaling in precision, trained and validated on the common crawl split of the Dolma dataset [Soldaini et al., 2024]. We use a standard causal Transformer++ implementation: SwiGLU activations [Shazeer, 2020], RoPE embeddings [Su et al., 2021], RMSLayerNorm, Adam $\beta$ values of $(0.9,0.95)$. We adopt a cosine learning rate schedule with 10% warmup period and peak learning rate of 6e-4 for the smallest model and learning rates scaled with width and depth according to depth-$\mu$P for the larger models [Yang et al., 2022, Bordelon et al., 2023]. We use a sequence length of 1024 and batch size of 256 throughout, with Adam $\epsilon$ 1e-15, following [Wortsman et al., 2023a]. We use weight decay of $0.1$, as [Ahmadian et al., 2023] find some results in the quantization literature may be artifacts of insufficient weight decay. We follow [Ma et al., 2024] in including a LayerNorm before projections because they find it is important for low precision training to be stable. These are the hyperparameters and settings used for the main scaling law experiments.
@@ Appendix B Additional Related Work
Low precision has been key to improving the efficiency of training and using LLMs [Micikevicius et al., 2017, Shoeybi et al., 2019, Wortsman et al., 2023b, Zhu et al., 2023]. Prior works generally study either precision during training [Courbariaux et al., 2014, Dettmers et al., 2024, 2021, Sun et al., 2020, Liu et al., 2023] or the effects of changing the precision after training (post-training quantization) [Frantar et al., 2022, Lin et al., 2024b, Dettmers et al., 2022, Xiao et al., 2023, Sheng et al., 2023, Dettmers et al., 2023]. In this work we study both, the precision during training and after, and unify them from a scaling perspective. Other important works include recent popular work on quantization-aware-training [Ma et al., 2024] where weights are quantized to extreme precisions (ternary) on the forward pass during training. This work is consistent with ours in that they can quantize weights so aggressively because weights are less sensitive than activations or KV cache. Further, while we use a fixed architecture throughout to maintain a controlled comparison across precision, they use a nonstandard architecture, learning rate, and weight decay schedule specifically to make training with ternary weights stable. Finally, [Keller, 2018] find similar tradeoffs for floating-point type with weights-only quantization.
@@
By scaling up the transformer architecture [Vaswani et al., 2017] a variety of large language models have been proposed [Brown et al., 2020, Rae et al., 2021, Touvron et al., 2023a, b, Dubey et al., 2024, Le Scao et al., 2023, Muennighoff et al., 2022, 2024a, Groeneveld et al., 2024, Jiang et al., 2023, Zhang et al., 2022, Allal et al., 2023, Li et al., 2023, Lozhkov et al., 2024, Luukkonen et al., 2023, Bai et al., 2023, Chowdhery et al., 2023, Team et al., 2023, Üstün et al., 2024, Deitke et al., 2024]. To improve our understanding of these models, various works have investigated their scaling properties [Ruan et al., 2024, Allen-Zhu and Li, 2024, Hägele et al., 2024]. Many aspects are relevant to scaling including the architecture [Tay et al., 2022a, Krajewski et al., 2024, Tao et al., 2024, Clark et al., 2022, Tay et al., 2022b, Scao et al., 2022, Peng et al., 2024], the modalities considered [Aghajanyan et al., 2023, Alabdulmohsin et al., 2022, Cherti et al., 2023], the performance metrics [Wei et al., 2022, Srivastava et al., 2022, Isik et al., 2024], the data composition [Li et al., 2024, Liu et al., 2024, Albalak et al., 2024] and data repetitions [Muennighoff et al., 2024b]. Our work analyzes one such aspect, which is key to better scaling: the numeric precision during and after training.